According to the Sleep Foundation, the average night’s sleep is 7 hours. Assume the standard deviation is 0.5 hours and that the probability distribution is normal.

a. What is the probability that a randomly selected person sleeps more than 9 hours?

b. What is the probability that a randomly selected person sleeps 5 hours or less?

c. What is the probability that a randomly selected person sleeps exactly 10 hours?

d. What is the probability that a randomly selected person sleeps more than 7 hours?

e. Doctors suggest getting between 7 and 9 hours of sleep each night. What percentage of the population gets this much sleep?
1
Expert's answer
2020-05-01T18:22:22-0400

a)P(x>9)


z=(x-m)/sd=(9-7)/0.5=4;


using table: P=1- t(z)=0.00003;


b) P(x<=5)


z=(5-7)/0.5=-4;


using table P=t(z)=0.00003 ;


c)P(X=10)=P(Z=((10-7)/0.5))=p(Z=6)=0


P=0%


d)P(x>7)


z=(7-7)/0.5=0;


using table P=1-t(0)=0.5


e)P(7<x<9)


z1=(7-7)/0.5=0;


z2=(9-7)/0.5=4;


using table P=t(4)-t(0)=0.99997-0.5=0.49997


P=49.997%